OpenStudy (anonymous):

The shorter leg of a 30°- 60°- 90° right triangle is 12.5 inches. How long are the longer leg and the hypotenuse?

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

In a 30-60-90 degree right triangle, the shorter leg is half of the hypotenuse, so the hypotenuse is 25. Now, let "x" be the length of the longer leg: x^2 + 12.5^2 = 25^2 x^2 = 25^2 - 12.5^2 So, just solve for "x" by squaring both terms on the right, adding them, and then taking the square root of that sum.

OpenStudy (anonymous):

x=21

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That's rounding a bit too much.

OpenStudy (anonymous):

wait i have to solve both ?

OpenStudy (anonymous):

approx. 21.65 is a better answer.

OpenStudy (anonymous):

both what?

OpenStudy (anonymous):

so the ansewer is 21.65

OpenStudy (anonymous):

longer leg approximation

OpenStudy (anonymous):

It's an irrational number, so if you want it in decimal form, it has to be approximated. The decimal points will go on forever otherwise.